OpenStudy (anonymous):

Which of the following is a correct equation for the line passing through the point (-2,1) and having slope m=1/2 Check all that apply A. y=1/2x + 2 B.y=-2x + 1/2 C. x-2y= -4 D. y-1= 1/2(x+2)

OpenStudy (x3_drummerchick):

y = mx + b 1 = -2 (1/2) + b solve for b ^ 1= (-1) + b 2 = b therefore... y = 1/2x + 2

OpenStudy (anonymous):

Is that only way and the only one?

OpenStudy (x3_drummerchick):

yes, it would be 'A'